A copper cycle experiment takes copper atoms through reactions that produce copper compounds and complexes one after the other, finally producing elemental copper. Copper atoms are conserved throughout the process. Given that a student begins with 10.55 mL of a 0.19 M Cu(NO3)2 solution, how much copper should be isolated at the end of the cycle?
Answer #1

Moles of Cu moles of cu (NO3)2 - volume x concentration of cu(No3)2 = 10. 5 5/1000 x 0-19 = 0-002 004 5 01 Mass of Cu = moles